Light traveling in glass n=3/2 is incident on-air n=1. What is the angle of incidence that causes the light to experience total internal reflection?


1
Expert's answer
2022-05-15T17:02:49-0400

The critical angle is given by

"\theta_c=\sin^{-1}\frac{n_2}{n_1}"

"\theta_c=\sin^{-1}\frac{1}{3/2}=41.8^\circ"

"\theta \geq\theta_c"